# Byam: Fixing Breaking Dependency Updates with Large Language Models

### v2

@@ 1 Introduction
Modern software development heavily relies on third-party software libraries to improve developers’ productivity and time-to-market. The set of libraries used by a client project is referred to as the dependencies of that project. Developers access a library’s functionality through its provided Application Programming Interfaces (APIs). These APIs evolve over time to introduce new functionality, fix bugs, or address security vulnerabilities (Lamothe et al., 2021). While such evolution is necessary, it often comes at a cost: updates may introduce breaking changes that prevent client projects from compiling or running  (Lamothe et al., 2021; Reyes et al., 2024b ; Reyes et al., 2024a ). Failures resulting from API updates are referred to as breaking dependency updates. When a new version of a dependency results in a breaking dependency update, the developers using the library (i.e., the client developers) have to figure out how to update their code to fix the errors they face. This is known to be a tedious and time consuming maintenance task (Zhang et al., 2022; Ochoa et al., 2022).
To address this problem, we propose Byam, a pipeline that leverages Large Language Models (LLMs) to update client code in response to breaking dependency updates (Vaswani, 2017; Chang et al., 2024). We focus on breaking dependency updates that lead to compilation errors, which are the most common types of causes for breakage (Reyes et al., 2024b ; Reyes et al., 2024a ). Byam analyzes compilation errors caused by a dependency update, formulates prompts that incorporate relevant build and dependency information, and queries an LLM to generate updated code. We assess the effectiveness of Byam using the BUMP dataset  (Reyes et al., 2024b ), which provides reproducible cases of breaking dependency updates in Java projects. Our experiments systematically evaluate different LLMs, prompt design strategies, and information granularity levels.
@@ 2 Background
Breaking dependency updates are common in software projects (Brito et al., 2018b ), and often discourage developers from upgrading their dependencies (Dietrich et al., 2019; Ochoa et al., 2022; Venturini et al., 2023). However, keeping dependencies up to date is crucial for security, as outdated libraries may contain vulnerabilities  (Brito et al., 2018b ; Xavier et al., 2017; Brito et al., 2020; Larios Vargas et al., 2020; Salza et al., 2018). A breaking dependency update occurs when a library update introduces incompatibilities that cause client code to fail. This can occur either when developers manually update dependencies or when they use automated tools such as Dependabot or Renovate, which submit pull requests for updates. Such breakages are often changes in method signatures, deprecations, or class removals can cascade into compilation errors across the client project. While breaking updates can also trigger test failures, dependency resolution issues or lock conflicts (Reyes et al., 2024b ), in this paper we focus specifically on compilation errors, which are the most prevalent form (Reyes et al., 2024a ).
@@
Prior work has extensively studied the prevalence and causes of breaking changes and methods to identify them in library code (Brito et al., 2018b ; Xavier et al., 2017; Brito et al., 2020; Brito et al., 2018a ; Mujahid et al., 2020; Reyes et al., 2024a ). Migration support has also been explored through rule-based transformations and recommendation systems (Xing and Stroulia, 2007; Dagenais and Robillard, 2009; Brito et al., 2018a ). However, such approaches typically rely on predefined rules, mined examples, or static analysis heuristics, which limit their ability to handle the diversity of real-world breaking updates and the project-specific context required to repair client code. In contrast, we investigate whether large language models (LLMs) can directly generate client-side repairs when provided with appropriate context.
@@
For systematic evaluation in this study, we examine breaking dependency updates as represented in the BUMP dataset Reyes et al., 2024b : a pair of commits for a project, consisting of a pre-breaking commit with a passing build and a breaking commit that updates the version of a single dependency, causing the build to fail.
@@ 3.3 Step 2: Use LLM to Generate Repairs
the error is triggered in the following specific lines in the previous code: <buggy_line>
@@
Providing API Difference Information (APIDiff) When a dependency update breaks a project, developers often compare the previous and new versions of the API to identify relevant changes, such as method renaming or modifications to the parameters (Brito et al., 2018a ). Such changes help developers determine how to update their client code to fix the problem. Following this practice, our intuition is that by explicitly providing these API differences (APIDiff) to the LLM, we can allow it to reason about the likely source of the error and suggest a more accurate solution.
To extract these differences, we rely on japicmp (Maven, 2022), a widely used tool to detect API changes in Java libraries. japicmp performs a comprehensive structural comparison, allowing it to detect modifications across public, internal, and even deprecated parts of an API. This tool identifies modifications such as changes in method signatures, parameter types, deprecated methods, or new functionality. For example, when we use japicmp to analyze the changes between versions 6.18.1 and 6.19.1 of the jasperreports dependency, it detects that the setLinewith method was removed from the new version of the dependency. We augment the prompt with that information, as in the example shown in Figure 6. Figure 7 shows the concrete example of APIDiff for jasperreports.
@@
Applying Chain of Thought (CoT) Prompting. Developers often solve complex errors by following a structured step-by-step reasoning process. This structured approach helps them to analyze the problem, identify possible causes and systematically find a solution. This structured approach is replicated in prompting in what is referred to as Chain of Thought (CoT) prompting (Wei et al., 2023). Previous work has shown that CoT instructions improve performance in code generation and automated repair tasks (Le Goues et al., 2019; Monperrus, 2018). Consequently, we consider CoT prompting for fixing breaking dependency updates. In our approach, CoT prompting explicitly guides the LLM to analyze the error message, identify the affected code, infer the underlying issue, and reason about possible fixes before proposing a corrected version. This approach guides the LLM to produce a more structured and coherent response (Wei et al., 2023). We incorporate the CoT technique in the prompt as shown in Figure 8.
@@
Prompt ID Prompt Name Client Code Error Message Buggy Line APIDiff CoT Prompting
@@ 4.2 Study Subjects
Our experiments are based on the real-world dependency updates collected in previous research Reyes et al., 2024b . We use BUMP, a benchmark for reproducible breaking dependency updates. BUMP includes 571 breaking dependency updates, of which 243 (43%), are broken builds due to compilation failure.
BUMP classifies compilation failures into four distinct categories: Direct compilation errors, Indirect compilation errors, Java version incompatibility and Werror failure (Reyes et al., 2024a ). We discard the 78 failures classified under the Java version incompatibility category. Java versioning failures are failures that require a different Java version than the one used in the project. Resolving Java versioning failures requires changing the Java version, rather than updating the code. BUMP also contains 8 compilation errors of type Werror failure, which is a failure due to activating the failOnWarning option in the configuration file after the dependency update. Since this type of failure relates to the linter configuration rather than the code, we disregard these data points. We discard cases in which, when resolving dependencies, the new version of the dependency conflicts with the dependencies declared in the Maven configuration file, causing errors when invoking API calls (Bono et al., 2024). These failures appear in the Direct compilation errors and Indirect compilation errors categories. Such cases are related to classpath issues, they are impossible to solve by updating the Java code. This leaves us with a total of 103 breaking dependency updates for our evaluation.
@@ 4.4 Language Models for Experimentation
We select the models we experiment with based on the results of RepairBench (Silva and Monperrus, 2024) and LiveCodeBench (Jain et al., 2024), which are established evaluation frameworks for LLMs in program repair. Based on these frameworks, we select the following five LLMs to use in our evaluation:
@@ 5 Experimental Results
Prompt
ID
Deepseek
V3
Gemini
2.0-flash
Gpt
4o-mini
o3
mini
Qwen2.5
32b-instruct
@@
Prompt
ID
Deepseek
V3
Gemini
2.0-flash
Gpt
4o-mini
o3
mini
Qwen2.5
32b-instruct
@@
Prompt
ID
Deepseek
V3
Gemini
2.0-flash
Gpt
4o-mini
o3
mini
Qwen2.5
32b-instruct
@@ 6.1 Comparison with Prior Work
In Frunkte and Krinke, (2025), the authors introduce two new approaches to fix breaking dependency updates using LLMs: a zero-shot prompting method and an agent system. Their agent iteratively attempts up to 30 fixes per case. The authors process the Bump dataset to produce two distinct evaluation data sets, the first set is considered as “light slice” with 65 projects with breaking dependency updates caused by only one file, and a “full slice” of 140 projects that has multiple files causing breaking in build. The authors test the zero-shot approach on the light-slice dataset while they test the agent system on both datasets. The zero-shot baseline uses no prompt engineering or design and is evaluated only on the light-slice dataset. The agent system, by contrast, is evaluated on both slices and operates iteratively, making up to 30 repair attempts per case. Importantly, the study does not report single-iteration success rates for the agent system, as its evaluation focuses on cumulative results across multiple trials.
Byam’s key advantage over the work in Frunkte and Krinke, (2025) is its context-rich prompting, evident in being able to achieve better performance using a less complicated approach . Byam fixed full builds that can consist of multiple file edits, achieving a success rate of 27% using the richest prompt (P8). Even when the full build does not succeed, Byam achieves a 78% fix rate of the individual compilation errors.
In contrast, the agent approach in Frunkte and Krinke, (2025) fixes a maximum of 23% of builds in the full slice and a maximum of 19% in the light-slice, with the zero-shot approach achieving a maximum success rate of 19%. We note that Frunkte and Krinke, (2025) does not report the average trials for fixes but allow iteration for up to 30 trials.
To provide a direct comparison of identical breaking dependency updates, we analyze the intersection of breaking dependency updates evaluated by both studies. The intersection between the 140 (“full slice”) breaking updates presented by Frunkte and Krinke, (2025) and the 103 breaking updates analyzed here yields 97 common breaking dependency updates. When evaluating Byam with o3-mini using the P8 configuration on these 97 shared cases, Byam achieves a 27% (26/97) success rate, compared to the agent approach’s 20% (19/97) performance on the same breaking dependency updates. This controlled comparison on the same failing builds eliminates dataset variability and demonstrates that our contextualized prompting strategy outperforms the agent-based approach.
@@ 8 Related Work
Breaking dependency updates is a common problem in software development. (Brito et al., 2018b ; Xavier et al., 2017; Brito et al., 2020) investigate the reasons behind the developer’s decision to introduce updates that lead to breaking dependency updates. These reasons include introducing new features, refactoring the code, fixing bugs, and addressing security vulnerabilities. Hejderup et al. analyze how changes in method behavior trigger incompatibilities despite preserving the API contract. Such changes are difficult to detect by static analysis and often only become apparent at runtime (Hejderup and Gousios, 2021). (Venturini et al., 2023) emphasized the challenges developers face in updating client code after dependency updates. Researchers investigated various approaches to address this issue, where some investigated detecting breaking dependency updates (Brito et al., 2018a ; Mujahid et al., 2020). (Reyes et al., 2024a ) recently introduced an automated approach to explain the breaking in the client code after the version update. Our approach differs from the previous studies as it uniquely addresses compilation failures caused by breaking updates through fixes generated using LLM. Byam originally exploits contextual information such as APIDiff and erroneous lines to augment prompts, a dimension unexplored in previous studies.
@@
Different approaches to mitigate breaking dependency updates are built on rule-based program transformation techniques. (Dagenais and Robillard, 2009) introduces a novel approach to API evolution by recommending adaptive replacements for deleted methods through analysis of repository history. (Xing and Stroulia, 2007) presents Diff-CatchUp, a tool that assists client application migration by automatically identifying broken APIs, suggesting plausible replacements, and providing usage examples based on model differencing and the framework’s actual working code. More recent research focuses on learning migration patterns from existing code samples (Lamothe et al., 2022), analyzing changes in customer projects that have already migrated (Xu et al., 2019), and leveraging documentation (Ni et al., 2021). These approaches are intended to reduce the manual effort required to update dependencies and increase the likelihood that developers maintain updated dependencies. Contrary to rule-based systems, our approach dynamically adapts to different types of breaking changes by integrating APIDiff and CoT reasoning. This eliminates the need for manual engineering of rules or migration examples by leveraging the generalization capabilities of LLMs.
@@
Large Language Models (LLMs) have emerged as a solution to automate API migration and dependency updates. Almeida et al. explored the use of ChatGPT for these tasks (Almeida et al., 2024). Compared to rule-based approaches, LLMs can handle contextual variations in the code, which makes them more effective in certain transformations (Nikolov et al., 2025). Tools based on LLMs have been developed to address specific migration problems. RELANCER employs machine learning to predict repairs on deprecated APIs in Jupyter notebooks (Zhu et al., 2021). PCART automates compatibility assessment and repair of API parameter changes in Python (Zhang et al., 2024). In addition, Liu et al. note that LLMs can repair tests affected by code changes (Liu et al., 2024).
@@
The most closely related work to our study is presented in parallel research by  (Frunkte and Krinke, 2025). The authors explore automated build repair using either zero-shot prompting or agent-based iterative repair loops without explicit prompt engineering, achieving at most 23% repair on the full-slice dataset. Byam improves on this by incorporating contextual information, reaching 27% success on the same dataset of builds in a single iteration and fixing 78% of individual compilation errors.

### v3

@@ 1 Introduction
Modern software development heavily relies on third-party software libraries to improve developers’ productivity and time-to-market. The set of libraries used by a client project is referred to as the dependencies of that project. Developers access a library’s functionality through its provided Application Programming Interfaces (APIs). These APIs evolve over time to introduce new functionality, fix bugs, or address security vulnerabilities (Lamothe et al. 2021). While such evolution is necessary, it often comes at a cost: updates may introduce breaking changes that prevent client projects from compiling or running  (Lamothe et al. 2021; Reyes et al. 2024b, a). Failures resulting from API updates are referred to as breaking dependency updates. When a new version of a dependency results in a breaking dependency update, the developers using the library (i.e., the client developers) have to figure out how to update their code to fix the errors they face. This is known to be a tedious and time consuming maintenance task (Zhang et al. 2022; Ochoa et al. 2022).
To address this problem, we propose Byam, a pipeline that leverages Large Language Models (LLMs) to update client code in response to breaking dependency updates (Vaswani 2017; Chang et al. 2024). We focus on breaking dependency updates that lead to compilation errors, which are the most common types of causes for breakage (Reyes et al. 2024b, a). Byam analyzes compilation errors caused by a dependency update, formulates prompts that incorporate relevant build and dependency information, and queries an LLM to generate updated code. We assess the effectiveness of Byam using the BUMP dataset  (Reyes et al. 2024b), which provides reproducible cases of breaking dependency updates in Java projects. Our experiments systematically evaluate different LLMs, prompt design strategies, and information granularity levels.
@@ 2 Background
Breaking dependency updates are common in software projects (Brito et al. 2018b), and often discourage developers from upgrading their dependencies (Dietrich et al. 2019; Ochoa et al. 2022; Venturini et al. 2023). However, keeping dependencies up to date is crucial for security, as outdated libraries may contain vulnerabilities  (Brito et al. 2018b; Xavier et al. 2017; Brito et al. 2020; Larios Vargas et al. 2020; Salza et al. 2018). A breaking dependency update occurs when a library update introduces incompatibilities that cause client code to fail. This can occur either when developers manually update dependencies or when they use automated tools such as Dependabot or Renovate, which submit pull requests for updates. Such breakages are often changes in method signatures, deprecations, or class removals can cascade into compilation errors across the client project. While breaking updates can also trigger test failures, dependency resolution issues or lock conflicts (Reyes et al. 2024b), in this paper we focus specifically on compilation errors, which are the most prevalent form (Reyes et al. 2024a).
@@
Prior work has extensively studied the prevalence and causes of breaking changes and methods to identify them in library code (Brito et al. 2018b; Xavier et al. 2017; Brito et al. 2020, 2018a; Mujahid et al. 2020; Reyes et al. 2024a). Migration support has also been explored through rule-based transformations and recommendation systems (Xing and Stroulia 2007; Dagenais and Robillard 2009; Brito et al. 2018a). However, such approaches typically rely on predefined rules, mined examples, or static analysis heuristics, which limit their ability to handle the diversity of real-world breaking updates and the project-specific context required to repair client code. In contrast, we investigate whether large language models (LLMs) can directly generate client-side repairs when provided with appropriate context.
@@
For systematic evaluation in this study, we examine breaking dependency updates as represented in the BUMP dataset Reyes et al. (2024b): a pair of commits for a project, consisting of a pre-breaking commit with a passing build and a breaking commit that updates the version of a single dependency, causing the build to fail.
@@ 3.3 Step 2: Use LLM to Generate Repairs
the error is triggered in the following specific lines in the previous code: <erroneous_line>
@@
Providing API Difference Information (APIDiff) When a dependency update breaks a project, developers often compare the previous and new versions of the API to identify relevant changes, such as method renaming or modifications to the parameters (Brito et al. 2018a). Such changes help developers determine how to update their client code to fix the problem. Following this practice, our intuition is that by explicitly providing these API differences (APIDiff) to the LLM, we can allow it to reason about the likely source of the error and suggest a more accurate solution.
To extract these differences, we rely on japicmp (Maven 2022), a widely used tool to detect API changes in Java libraries. japicmp performs a comprehensive structural comparison, allowing it to detect modifications across public, internal, and even deprecated parts of an API. This tool identifies modifications such as changes in method signatures, parameter types, deprecated methods, or new functionality. For example, when we use japicmp to analyze the changes between versions 6.18.1 and 6.19.1 of the jasperreports dependency, it detects that the setLinewith method was removed from the new version of the dependency. We augment the prompt with that information, as in the example shown in Figure 6. Figure 7 shows the concrete example of APIDiff for jasperreports.
@@
Applying Chain of Thought (CoT) Prompting. Developers often solve complex errors by following a structured step-by-step reasoning process. This structured approach helps them to analyze the problem, identify possible causes and systematically find a solution. This structured approach is replicated in prompting in what is referred to as Chain of Thought (CoT) prompting (Wei et al. 2023). Previous work has shown that CoT instructions improve performance in code generation and automated repair tasks (Le Goues et al. 2019; Monperrus 2018). Consequently, we consider CoT prompting for fixing breaking dependency updates. In our approach, CoT prompting explicitly guides the LLM to analyze the error message, identify the affected code, infer the underlying issue, and reason about possible fixes before proposing a corrected version. This approach guides the LLM to produce a more structured and coherent response (Wei et al. 2023). We incorporate the CoT technique in the prompt as shown in Figure 8.
@@
Prompt ID Prompt Name Client Code Error Message Erroneous Line APIDiff CoT Prompting
@@ 4.2 Study Subjects
Our experiments are based on the real-world dependency updates collected in previous research Reyes et al. (2024b). We use BUMP, a benchmark for reproducible breaking dependency updates. BUMP includes 571 breaking dependency updates, of which 243 (43%), are broken builds due to compilation failure.
BUMP classifies compilation failures into four distinct categories: Direct compilation errors, Indirect compilation errors, Java version incompatibility and Werror failure (Reyes et al. 2024a). We discard the 78 failures classified under the Java version incompatibility category. Java versioning failures are failures that require a different Java version than the one used in the project. Resolving Java versioning failures requires changing the Java version, rather than updating the code. BUMP also contains 8 compilation errors of type Werror failure, which is a failure due to activating the failOnWarning option in the configuration file after the dependency update. Since this type of failure relates to the linter configuration rather than the code, we disregard these data points. We discard cases in which, when resolving dependencies, the new version of the dependency conflicts with the dependencies declared in the Maven configuration file, causing errors when invoking API calls (Bono et al. 2024). These failures appear in the Direct compilation errors and Indirect compilation errors categories. Such cases are related to classpath issues, they are impossible to solve by updating the Java code. This leaves us with a total of 103 breaking dependency updates for our evaluation.
@@ 4.4 Language Models for Experimentation
We select the models we experiment with based on the results of RepairBench (Silva and Monperrus 2024) and LiveCodeBench (Jain et al. 2024), which are established evaluation frameworks for LLMs in program repair. Based on these frameworks, we select the following five LLMs to use in our evaluation:
@@ 5 Experimental Results
Prompt ID Deepseek V3 Gemini 2.0-flash Gpt 4o-mini o3 mini Qwen2.5 32b-instruct
@@
Prompt ID Deepseek V3 Gemini 2.0-flash Gpt 4o-mini o3 mini Qwen2.5 32b-instruct
@@
Prompt ID Deepseek V3 Gemini 2.0-flash Gpt 4o-mini o3 mini Qwen2.5 32b-instruct
@@ 6.1 Comparison with Prior Work
In Frunkte and Krinke (2025), the authors introduce two new approaches to fix breaking dependency updates using LLMs: a zero-shot prompting method and an agent system. Their agent iteratively attempts up to 30 fixes per case. The authors process the Bump dataset to produce two distinct evaluation data sets, the first set is considered as “light slice” with 65 projects with breaking dependency updates caused by only one file, and a “full slice” of 140 projects that has multiple files causing breaking in build. The authors test the zero-shot approach on the light-slice dataset while they test the agent system on both datasets. The zero-shot baseline uses no prompt engineering or design and is evaluated only on the light-slice dataset. The agent system, by contrast, is evaluated on both slices and operates iteratively, making up to 30 repair attempts per case. Importantly, the study does not report single-iteration success rates for the agent system, as its evaluation focuses on cumulative results across multiple trials.
Byam’s key advantage over the work in Frunkte and Krinke (2025) is its context-rich prompting, evident in being able to achieve better performance using a less complicated approach . Byam fixed full builds that can consist of multiple file edits, achieving a success rate of 27% using the richest prompt (P8). Even when the full build does not succeed, Byam achieves a 78% fix rate of the individual compilation errors.
In contrast, the agent approach in Frunkte and Krinke (2025) fixes a maximum of 23% of builds in the full slice and a maximum of 19% in the light-slice, with the zero-shot approach achieving a maximum success rate of 19%. We note that Frunkte and Krinke (2025) does not report the average trials for fixes but allow iteration for up to 30 trials.
To provide a direct comparison of identical breaking dependency updates, we analyze the intersection of breaking dependency updates evaluated by both studies. The intersection between the 140 (“full slice”) breaking updates presented by Frunkte and Krinke (2025) and the 103 breaking updates analyzed here yields 97 common breaking dependency updates. When evaluating Byam with o3-mini using the P8 configuration on these 97 shared cases, Byam achieves a 27% (26/97) success rate, compared to the agent approach’s 20% (19/97) performance on the same breaking dependency updates. This controlled comparison on the same failing builds eliminates dataset variability and demonstrates that our contextualized prompting strategy outperforms the agent-based approach.
@@ 6.2 Unsuccessful Repairs and Test Failures
6.3 Recent Related Industry Tools
Over the past year, in parallel with this research, several industrial tools have emerged that use large language models and agents to support software development tasks, including dependency updates and related software maintenance tasks. Some of these tools are “general-purpose development agents” while some specifically focus on dependency updates.
General-purpose coding agents that can be assigned various tasks across the software development process have become increasingly available. Such agents can be standalone CLI tools (e.g., Co-pilot CLI  [^7] https://github.com/features/copilot/cli or Gemini CLI  [^8] https://geminicli.com/) or can be directly integrated into development platforms such as GitHub (e.g. GitHub Copilot  [^9] https://github.com/features/copilot) or Integrated Development Environments (IDEs) such as VSCode (e.g., VSCode Copilot  [^10] https://github.com/features/copilot/ai-code-editor or Cursor  [^11] https://cursor.com/agents). Once prompted with a task, the agents can run in the background to complete the assigned task. These agents can be assigned issues to solve, whether the issue is about fixing a bug or implementing a new feature. For GitHub-integrated tools, the agent may create a PR with the required changes based on the prompt or assigned issue. If it is integrated into an IDE, it can start generating or editing files directly in the editor. While these agents do not specifically focus on fixing breaking dependency updates, the task can be viewed as a specific instance of program repair that these agents can handle. Generally speaking, these industrial agents are designed as end-to-end automation pipelines that integrate multiple stages of repair, such as fault diagnosis, patch generation, testing, and deployment—within real development workflows. Accordingly, while developers can use them to fix breaking dependency updates, we do not have empirical evidence about their effectiveness for this specific task.
On the other hand, there are some recent industrial tools that are more closely related to breaking dependency updates. For example, Dynatrace  [^12] https://www.dynatrace.com/ and Develocity  [^13] https://docs.gradle.com/develocity/current/ provide in-depth monitoring of build status. These tools analyze build and test data, with the aim of providing insights into aspects like performance, reliability, and efficiency. While these tools do not attempt to fix the breaking update themselves, they can potentially provide richer context that can help the LLM provide a correct fix. Experimenting with the effect of this additional context is an interesting avenue for future work.
The closest to our setting are tools that specialize in fixing breaking dependency updates. For example, FOSSA  [^14] https://fossa.com/products/fossabot/ automatically reviews updates for breaking changes and analyzes the code impact of updates, working alongside tools like Dependabot and Renovate to help developers manage dependencies. Patchwork  [^15] https://docs.patched.codes/patchwork/overview is another example. These tools support automated repair by performing steps such as comparing dependency versions, generating patches, validating fixes, and potentially opening pull requests with proposed changes. However, similar to general-purpose agents, they are designed as end-to-end automation pipelines that integrate multiple stages of repair, rather than enabling in-depth experimentation and analysis of individual steps. In contrast, our study evaluates LLM-based repair under scientifically controlled conditions. We focus on empirical evidence for specific repair principles, including (i) the role of targeted contextual information about the breaking change, (ii) fault localization that highlights failing lines and relevant API changes, and (iii) prompt design choices and how they affect repair outcomes. Our goal is to isolate these factors and analyze how they influence repair effectiveness for breaking changes. Thus, rather than providing a full end-to-end automation solution, our work studies the repair step in isolation to understand which design choices contribute to successful outcomes. Our results can inform improvements to relevant academic and industrial tools.
6.4 Industry versus Academia in Software Research
The previous section demonstrates that since we began this research, the broader AI-assisted software development ecosystem has advanced quickly, especially with advances in the capabilities of large language models and agent-based systems. This forces us to reflect on these pace differences and their impact.
Generally speaking, industrial progress is fast paced with rapid updates in response to user feedback and needs. On the other hand, the requirements of academic research are different. It requires careful experimental design, dataset construction, controlled evaluation, peer review, and reproducibility. In other words, industrial tools drive adoption and deliver end-to-end automation, but they are typically evaluated through operational metrics and real-world usage rather than controlled experimental measurements of individual design choices. As a result, it remains difficult to assess how well their behavior generalizes or to reason about the causes of success and failure. Research, on the other hand, systematically investigates techniques and provides empirical evidence about what works, under which conditions, and why. These differing objectives naturally lead to different evaluation practices and timelines between industrial development and academic research.
@@ 8 Related Work
Breaking dependency updates is a common problem in software development. (Brito et al. 2018b; Xavier et al. 2017; Brito et al. 2020) investigate the reasons behind the developer’s decision to introduce updates that lead to breaking dependency updates. These reasons include introducing new features, refactoring the code, fixing bugs, and addressing security vulnerabilities. Hejderup et al. analyze how changes in method behavior trigger incompatibilities despite preserving the API contract. Such changes are difficult to detect by static analysis and often only become apparent at runtime (Hejderup and Gousios 2021). (Venturini et al. 2023) emphasized the challenges developers face in updating client code after dependency updates. Researchers investigated various approaches to address this issue, where some investigated detecting breaking dependency updates (Brito et al. 2018a; Mujahid et al. 2020). (Reyes et al. 2024a) recently introduced an automated approach to explain the breaking in the client code after the version update. Our approach differs from the previous studies as it uniquely addresses compilation failures caused by breaking updates through fixes generated using LLM. Byam originally exploits contextual information such as APIDiff and erroneous lines to augment prompts, a dimension unexplored in previous studies.
@@
Different approaches to mitigate breaking dependency updates are built on rule-based program transformation techniques. (Dagenais and Robillard 2009) introduces a novel approach to API evolution by recommending adaptive replacements for deleted methods through analysis of repository history. (Xing and Stroulia 2007) presents Diff-CatchUp, a tool that assists client application migration by automatically identifying broken APIs, suggesting plausible replacements, and providing usage examples based on model differencing and the framework’s actual working code. More recent research focuses on learning migration patterns from existing code samples (Lamothe et al. 2022), analyzing changes in customer projects that have already migrated (Xu et al. 2019), and leveraging documentation (Ni et al. 2021). These approaches are intended to reduce the manual effort required to update dependencies and increase the likelihood that developers maintain updated dependencies. Contrary to rule-based systems, our approach dynamically adapts to different types of breaking changes by integrating APIDiff and CoT reasoning. This eliminates the need for manual engineering of rules or migration examples by leveraging the generalization capabilities of LLMs.
@@
Large Language Models (LLMs) have emerged as a solution to automate API migration and dependency updates. Almeida et al. explored the use of ChatGPT for these tasks (Almeida et al. 2024). Compared to rule-based approaches, LLMs can handle contextual variations in the code, which makes them more effective in certain transformations (Nikolov et al. 2025). Tools based on LLMs have been developed to address specific migration problems. RELANCER employs machine learning to predict repairs on deprecated APIs in Jupyter notebooks (Zhu et al. 2021). PCART automates compatibility assessment and repair of API parameter changes in Python (Zhang et al. 2024). In addition, Liu et al. note that LLMs can repair tests affected by code changes (Liu et al. 2024).
@@
The most closely related work to our study is presented in parallel research by  (Frunkte and Krinke 2025). The authors explore automated build repair using either zero-shot prompting or agent-based iterative repair loops without explicit prompt engineering, achieving at most 23% repair on the full-slice dataset. Byam improves on this by incorporating contextual information, reaching 27% success on the same dataset of builds in a single iteration and fixing 78% of individual compilation errors.